OpenStudy (flix):

find limit lim┬((x,y)→(0,0)) (5x^2+5y^2)/√(x^2+y^2 ).

OpenStudy (anonymous):

\[5\frac{ x^2 + y^2 }{ \sqrt{x^2 + y^2} } = 5\sqrt{x^2 + y^2}\] plug in (x, y) = (0, 0) answer is 0

OpenStudy (flix):

thank you so much..